Exercise 5.84 Calculate the root mean square velocity and kinetic energy of CO, CO2, and SO3 at 296 K. Rev Part A Calculate t circle answers
Answer #1

A)

Molar mass of CO,

mm = 28.01 g/mol

= 0.02801 Kg/mol

T = 296 K

vrms = sqrt (3*R*T/M)

here : T is temperature in Kelvin

and M is molar mass in Kg/mol

So,

vrms = sqrt(3*8.314*296.0/0.02801)

= 5.134*10^2 m/s

Answer: 5.13*10^2 m/s

B)

Molar mass of CO2,

mm = 44.01g/mol

= 0.04401 Kg/mol

T = 296 K

vrms = sqrt (3*R*T/M)

here : T is temperature in Kelvin

and M is molar mass in Kg/mol

So,

vrms = sqrt(3*8.314*296.0/0.04401)

= 4.096*10^2 m/s

Answer: 4.10*10^2 m/s